​Pete’s Tennis Shop has the following transactions related to its top-selling Wilson tennis racket for the month of August. Pete’s Tennis Shop uses a periodic inventory system:

Cost per Unit Date August 1 August 4 August 11 August 13 August 20 August 26 August 29 Transactions Units Total Cost $141 $ 1,128 1,310 1,210 1,221 $ 4,869 Beginning inventory Sale ($130 each) Purchase Sale ($145 each) 8 5 10 8 10 131 Prchase 121 Sale ($155 each) Purchase

1. Calculate ending inventory and cost of goods sold at August 31, using the specific identification method.

(For the specific identification method, the August 4 sale consists of rackets from beginning inventory, the. August 13 sale consists of rackets from the August 11 purchase, and the August 26 sale consists of one racket from beginning inventory and 10 rackets from the August 20 purchase.)
2. Using FIFO, calculate ending inventory and cost of goods sold at August 31.

3. Using LIFO, calculate ending inventory and cost of goods sold at August 31.  

4. Using weighted-average cost, calculate ending inventory and cost of goods sold at August 31. (Round intermediate and final answers to 2 decimal places).

5. Calculate sales revenue and gross profit under each of the four methods. (Round eighted-average cost amounts to 2 decimal places)

6. Comparing FIFO and LIFO, which one provides the more meaningful measure of ending inventory? (FIFO or LIFO)

7. If Pete’s chooses to report inventory using LIFO, record the LIFO adjusting journal entry for Augist 31.  
Answer #1

Calculate the requirements as follows: 1 Total units 2 Units sold 3 Units in ending inventory 4 Weighted average cost per uniCalculate value in above table as follows: в A 1 Total units =8+10+10+11 2 Units sold =5+8+11 I=B1-B2 4 Weighted average cost5. Calculate sales revenue and gross profit under each of the four methods as follows: B A 1 Particulars 2 Sales Revenue (A)FIFO method provides more meaningful measure of ending inventory as it provides the inventory which is most recently purchase7. Record LIFO adjustment entry as follows: ДА В 1 Date Account Title & Explanation 2 31-Aug Inventory Cost of Goods Sold C D